# Rule-based Data Selection for Large Language Models

### v1

@@ 1 Introduction
Recent studies now adopt an approach that employs LLMs to grade data quality according to a set of designed metrics (which we call rules) (Yuan et al., 2024; Wettig et al., 2024; Bai et al., 2022; Mu et al., ). For example, Wettig et al. (2024) rates the pre-training data using LLMs according to four predefined rules. RedPajama (Together AI, 2023) is continuously developing a pool of rules for users to select from, which currently contains over 40 basic criteria that LLM data should satisfy. On specific aspects such as safety, Constitutional AI (Bai et al., 2022) proposed their “constitution”—a set of standard safety criteria—to generate safe synthetic data, and most recently, OpenAI’s Rule-based Rewarding (Mu et al., ) proposed 21 general safety rules and injected them into the RLHF (reinforcement learning with human feedback) process. This rule-based rating provides greater explainability of data quality and breaks down the challenge of assigning a data point one overall quality score into a simpler task of giving several rule-specific scores. Evidence suggests that this fine-grained approach also yields more accurate rating outcomes (Yuan et al., 2024; Wettig et al., 2024; Bai et al., 2022; Mu et al., ).
@@
Note that our data selection methodology is highly versatile and applicable to a variety of scenarios, including LLM pre-training, fine-tuning on specific domains, RLHF preference data, etc. We conduct experiments to cover a range of tasks and datasets, including general pre-training data and domain-specific data in four domains: IMDB, Medical, Math, and Code. We show that our rule-based data selection typically yields more accurate rating results, thereby enhancing data quality and leading to better performance of the LLM trained with the data. Here is a list of the main contributions of our work:
@@ 2 Related Work
Rule-based rating. There are studies adopting a more fine-grained approach to data quality, distilling it into a finite set of metrics which we refer to as “rules”. For instance, RedPajama (Together AI, 2023) provides over 40 quality rules that serve as basic quality metrics for the users to choose from. More pertinent to our research, there are papers that apply this rule-based idea to rate LLM data. For example, Yuan et al. (2024) assigns a score out of 5 to each data point, awarding 1 point for each of the 5 predefined criteria met. In Wettig et al. (2024), the authors designed four general rules to rate and select data for LLM pre-training. (Sun et al., 2024) proposed 16 human-crafted rules to evaluate the desirable quality of response data. The rule-based approach is also utilized in more targeted applications, such as ensuring data safety. Constitutional AI designed 16 general safety critique rules to revise synthetic data, enhancing data safety (Bai et al., 2022). This revision process involves iterative steps where a random subset of rules from the “constitution” (the entire set of rules) is applied. Additionally, in (Mu et al., ), the score generated by an LLM grader according to a set of 21 safety rules is integrated directly into the RLHF process as an additional reward. As noted earlier in the introduction, the rules employed in the literature exhibit several critical issues. They often depend heavily on human heuristics for design, lack robust rule evaluation metrics and exploration of rule sizes, and demonstrate limited versatility for new tasks or for customization. Our goal is to address these challenges using our proposed framework.
@@ 3.1 Definitions and Notations
Measure orthogonality: We propose a metric for selecting rules based on the orthogonality of score vectors. Here we introduce a mathematical definition to quantify the orthogonality or correlation of a set of score vectors. Given a score matrix $\bar{{\bm{S}}}\in\mathbb{R}^{n\times r}$ such that the columns are the score vectors of dimensions $n$. We begin by computing the sample covariance matrix $Cov(\bar{{\bm{S}}})$ for the columns of $\bar{{\bm{S}}}$, whose entries are defined by
$Cov(\bar{{\bm{S}}})_{i,j}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{n-1}\sum_{k=1}^{n}(S_{k,i}-\mu_{i})(S_{k,j}-\mu_{j}),\qquad 1\leq i,j\leq r\text{ and }i\neq j,$
$Cov(\bar{{\bm{S}}})_{i,i}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{n-1}\sum_{k=1}^{n}(S_{k,i}-\mu_{i})^{2},\qquad 1\leq i\leq r,$
@@ 4.2 Results
DPP rules v.s. Randomly selected rules (answer to Q4). We compare DPP-selected rules with randomly selected rules of the same size $r$, evaluating both the rule correlation $\rho(\bar{{\bm{S}}})$ and the MSE $\epsilon(\bar{{\bm{S}}})$ for their corresponding score submatrices $\bar{{\bm{S}}}$. The results show that DPP consistently produces rules with lower correlation and lower MSE, regardless of the value of $r$ (see Figures 3(b) and 3(c)). Another key observation is that the MSE for DPP rules increases when $r$ is either too small or too large, with the optimal $r$ falling somewhere in the middle. This matches our intuition: when $r$ is too small, there are too few rules to achieve sufficient rating diversity, and when $r$ is too large, rule redundancy can negatively affect the rating outcomes.
@@ 5 Evaluation B: Data Selection for LLM Fine-tuning
In this section, we follow the pipeline outlined in Section 3.3 and conduct experiments based on Evaluation B, where we train an LLM (Llama3-8B) using the selected data and assess its performance. This setup closely reflects real-world applications of LLM data selection. We benchmark our method against several baselines, including random selection, direct rating without rules, QuRating rules (Wettig et al., 2024), and DSIR (Xie et al., 2024) (a commonly used baseline for LLM data selection). We condcut experiments in this section to explore the following research questions: (Q1) How does data selected by rule-based methods enhance model fine-tuning compared to rule-free methods? (Q2) How do the rules generated by our automated framework compare to human-designed rules? (Q3) How does DPP rule selection perform compared to random rule selection?
@@ 5.1 Experiments Setup
Compared Methods. We compare our method against the following baselines, where the first three represent rule-free data selection methods, and QuRating represents a rule-based method using human-designed rules.
  • Uniform sampling: A standard baseline where data is uniformly sampled from the source. The results are averaged over 2 independent trials.
  • No Rule: We prompt Llama3-8B (instruction-tuned version) to individually rate the data without rules, and then apply the same sampling procedure as described in Step 4 of 3.3.
  • DSIR (Xie et al., 2024): A framework that uses importance resampling to sample data that resemble a target dataset (we use Wikipedia as the target for the general continued pre-training and used benchmark test datasets for the domain fine-tuning) .
  • QuRating (Wettig et al., 2024): Data are rated based on four rules: writing style, required expertise, facts & trivia, and educational value. We only implement QuRating in the general continued pre-training experiments, as these qualities are not designed for any specific domain.
For our automated rule-based selection algorithm, we set $R=50$, as in Section 4, and select $r=10$ for rule selection. As inferences of LLM are a lot less computing-consuming than model training, we set $n=10^{4}$ in all our experiments. The choice of $r$ as a hyperparameter is based on experimental observations from Section 4, where very small or very large values of $r$ did not yield optimal results. Discussion and exploration of different values for $r$ are provided in Appendix A.6.3, and details of the rule generation prompts, rating prompts, generated and selected rules are in Appendix A.6.7. To demonstrate the effectiveness of our automated rule-based selection algorithm, we evaluate the performance of the following methods developed within our framework, and we also conduct comparisons among these methods to demonstrate the advantages of DPP in rule selection.
@@
  • Random 10 Rules: Randomly choose 10 rules and average the score vectors to rate and select data. The results are averaged over 2 independent trials.
  • DPP 10 Rules: Use DPP to choose 10 rules, and then apply them to rate and select data.
@@ 5.2 General Continued Pre-training
We selected 20K samples from our data source using the methods described above for continued pre-training of Pythia-1B, then benchmarked the model’s performance. The choice of 20K samples (approximately 20M tokens) was constrained by our GPU resources. Although 20K is significantly smaller than the pre-training corpus size, we observed improvements in benchmark results shown in Table 1, with DPP consistently leading in most metrics.
@@
All 50 Rules 60.0 25.3 53.0 26.1 49.3 42.7
Random 10 Rules 60.0 25.1 53.7 25.8 49.5 42.8
DPP 10 Rules 60.0 25.7 54.0 26.2 50.1 43.2
Table 1: General continued pre-training of Pythia-1B using 20K selected data samples from SlimPajama. The first row shows the original model’s performance without further training.
@@ 5.3 Domain-specifc Fine-tuning
We now focus on domain-specific fine-tuning across four domains: IMDB, Medical, Math, and Code. By selecting 20K domain-related data samples from our source for model training, we aim to enhance domain-specific task performance. As demonstrated in Tables 2 and 3, domain-specific fine-tuning yields more significant improvements than general continued pre-training, which often needs larger datasets to enhance performance due to the broader nature of the training data. Notably, rule-based methods consistently outperform rule-free approaches in general, especially when comparing against Random Select and No Rule. Within the rule-based methods, DPP demonstrated superior performance compared to using all 50 rules or selecting 10 rules randomly. This underscores the effectiveness of a rule-based strategy, which introduces more diversity in the data rating step and selects better training data. Furthermore, it also demonstrates that our application of DPP in rule selection effectively identifies a core set of high-quality rules, thereby enhancing data quality and ultimately improving model performance.
@@
All 50 Rules 51.0 23.1 42.6 23.0 29.6
Random 10 Rules 51.7 24.0 41.2 23.5 29.6
DPP 10 Rules 53.3 24.3 43.0 26.0 31.1
@@
Llama3-8B 41.0 39.6 34.0 38.2 46.3 42.9 44.0 48.4 45.4
Random Select 40.5 39.2 35 38.2 38.7 38.2 38.2 39.7 38.7
No Rule 42.3 37.4 37.0 38.9 45.1 43.9 42.8 52.1 45.9
DSIR 41.5 41.1 34.0 38.9 45.1 43.6 49.1 52.2 47.5
All 50 Rules 41.8 40.7 33.0 38.5 43.9 43.4 46.6 49.1 45.8
Random 10 Rules 42.9 39.6 35.0 39.2 48.5 40.8 46.6 48.1 46.0
DPP 10 Rules 43.6 40.4 38.0 40.7 50.6 44.1 46.9 52.8 48.6
@@ A.2 DPP Sampling
Time Complexity Finding the submatrix (subset of columns) of a matrix to maximize the orthogonality is NP-hard (Civril & Magdon-Ismail, 2007; Kulesza et al., 2012). DPP provides us a relatively good solution. In practice, the computational complexity of sampling from a DPP depends primarily on the eigendecomposition of the kernel matrix ${\bm{K}}$. In our case, ${\bm{K}}\in\mathbb{R}^{R\times R}$ and therefore it requires $O(R^{3})$ time, where $R$ is the number of rules. In the DPPy package (Gautier et al., 2019) it uses the spectral sampler by default, so the actual run-time of our DPP implementation is $O(R^{3})$.
@@ A.5.4 Rater Model Size Variation: Llama3-70B-Instruct
A.5.5 Prompts and Generated Rules

### v2

@@ 1 Introduction
Recent studies now adopt an approach that employs LLM-as-a-judge to grade data quality according to a set of designed metrics (which we call rules) (Yuan et al., 2024; Wettig et al., 2024; Bai et al., 2022; Mu et al., ). For example, Wettig et al. (2024) rates the pre-training data using LLMs according to four predefined rules. RedPajama (Together AI, 2023) is continuously developing a pool of rules for users to select from, which currently contains over 40 basic criteria that LLM data should satisfy. On specific aspects such as safety, Constitutional AI (Bai et al., 2022) proposed their “constitution”—a set of standard safety criteria—to generate safe synthetic data, and in Huang et al. (2024) they have developed 133 rules. Most recently, OpenAI’s Rule-based Rewarding (Mu et al., ) proposed 21 general safety rules and injected them into the RLHF (reinforcement learning with human feedback) process. This rule-based rating provides greater explainability of data quality and breaks down the challenge of assigning a data point one overall quality score into a simpler task of giving several rule-specific scores. Evidence suggests that this fine-grained approach also yields more accurate rating outcomes (Yuan et al., 2024; Wettig et al., 2024; Bai et al., 2022; Mu et al., ).
@@
Note that our data selection methodology is highly versatile and applicable to a variety of scenarios, including LLM pre-training, fine-tuning on specific domains, RLHF preference data, etc. We conduct experiments to cover a range of tasks and datasets, including general pre-training data and domain-specific data in four domains: IMDB, Medical, Math, and Code. [^1] Code of our experiments: https://anonymous.4open.science/r/DataSelection-F118/ We show that our rule-based data selection typically yields more accurate rating results, thereby enhancing data quality and leading to better performance of the LLM trained with the data. Here is a list of the main contributions of our work:
@@ 2 Related Work
Rule-based rating. There are studies adopting a more fine-grained approach to data quality, distilling it into a finite set of metrics which we refer to as “rules”. For instance, RedPajama (Together AI, 2023) provides over 40 quality rules that serve as basic quality metrics for the users to choose from. More pertinent to our research, there are papers that apply this rule-based idea to rate LLM data. For example, Yuan et al. (2024) assigns a score out of 5 to each data point, awarding 1 point for each of the 5 predefined criteria met. In Wettig et al. (2024), the authors designed four general rules to rate and select data for LLM pre-training. (Sun et al., 2024) proposed 16 human-crafted rules to evaluate the desirable quality of response data. The rule-based approach is also utilized in more targeted applications, such as ensuring data safety. Constitutional AI designed 16 general safety critique rules to revise synthetic data, enhancing data safety (Bai et al., 2022). This revision process involves iterative steps where a random subset of rules from the “constitution” (the entire set of rules) is applied. Additionally, in Mu et al. , the score generated by an LLM grader according to a set of 21 safety rules is integrated directly into the RLHF process as an additional reward. In Wang et al. (2024), they design a composite reward model in RLHF, trained using rule-based ratings. As noted earlier in the introduction, the rules employed in the literature exhibit several critical issues. They often depend heavily on human heuristics for design, lack robust rule evaluation metrics and exploration of rule sizes, and demonstrate limited versatility for new tasks or for customization. Our goal is to address these challenges using our proposed framework.
@@ 3.1 Definitions and Notations
Measure orthogonality: We propose a metric for selecting rules based on the orthogonality of score vectors. Here we introduce a mathematical definition to quantify the orthogonality or correlation of a set of score vectors. Given a score matrix $\bar{{\bm{S}}}\in\mathbb{R}^{n\times r}$ such that the columns are the score vectors of dimensions $n$. We begin by computing the covariance matrix $Cov(\bar{{\bm{S}}})$ for the columns of $\bar{{\bm{S}}}$, whose entries are defined by
$Cov(\bar{{\bm{S}}})_{i,j}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{n}\sum_{k=1}^{n}(S_{k,i}-\mu_{i})(S_{k,j}-\mu_{j}),\qquad 1\leq i,j\leq r,$
@@ 4.2 Results
DPP rules v.s. Randomly selected rules (answer to Q4). We compare DPP-selected rules with randomly selected rules of the same size $r$, evaluating both the rule correlation $\rho(\bar{{\bm{S}}})$ and MSE $\epsilon(\bar{{\bm{S}}})$ for their corresponding score submatrices $\bar{{\bm{S}}}$. The results show that DPP consistently produces rules with lower correlation and MSE, regardless of the value $r$ (see Figures 3(b) and 3(c)). Another key observation is that the MSE for DPP rules increases when $r$ is either too small or too large, with the optimal $r$ falling somewhere in the middle. This matches our intuition: when $r$ is too small, there are too few rules to achieve sufficient rating diversity, and when $r$ is too large, rule redundancy can negatively affect the rating outcomes. In fact, this motivates our selection of $r=10$ in this paper.
@@ 5 Evaluation B: Data Selection for LLM Fine-tuning
In this section, we follow the pipeline outlined in Section 3.3 and conduct experiments based on Evaluation B, where we train an LLM (Llama3-8B) using the selected data and assess its performance. This setup closely reflects real-world applications of LLM data selection. We benchmark our method against several baselines, such as uniform sampling, direct rating without rules, QuRating rules (Wettig et al., 2024), and DSIR (Xie et al., 2024) (a commonly used baseline for LLM data selection). We condcut experiments in this section to explore the following research questions: (Q1) How does data selected by rule-based methods enhance model fine-tuning compared to rule-free methods? (Q2) How do the rules generated by our automated framework compare to human-designed rules? (Q3) How does DPP rule selection perform compared to random rule selection?
@@ 5.1 Experiments Setup
Compared Methods. We compare our method against the following baselines, including both rule-free and rule-based data selection methods. For rule-free methods, we have: Uniform sampling: select the data randomly, No Rule: prompt Llama3-8B-Instruction to individually rate the data without rules, and then apply the same sampling procedure as described in 3.3, DSIR (Xie et al., 2024): importance resampling of data that resemble a target dataset (we use Wikipedia as the target for the general continued pre-training and benchmark test datasets for the domain fine-tuning). For rule-based methods, we include: QuRating (Wettig et al., 2024): data rating and selection using four human-designed rules, and GPT-Uncorrelated: Directly prompting GPT-4 to generate 10 uncorrelated rules for data rating and selection. We have comparison against more baseline methods in Appendix A.6.2.
For our automated rule-based selection algorithm, we set $R=50$, as in Section 4, and select $r=10$ for rule selection. As inferences of LLM are a lot less computing-consuming than model training, we set $n=10^{4}$ in all our experiments. The choice of $r$ as a hyperparameter is based on experimental observations from Section 4, where very small or very large values of $r$ did not yield optimal results. Discussion and exploration of different values for $r$ are provided in Appendix A.6.5, and details of the rule generation prompts, rating prompts, generated and selected rules are in Appendix A.6.11. To demonstrate the effectiveness of our automated orthogonal-rule-based selection algorithm, we evaluate the performance of the following methods developed within our framework, and we also conduct comparisons among these methods to demonstrate the advantages of DPP in rule selection.
@@
  • Random 10 Rules: Randomly choose 10 rules and average the score vectors to rate and select data.
  • DPP 10 Rules: Use DPP to sample 10 rules, and then apply them to rate and select data.
Note that for uniform sampling and the methods involving randomness in the selection of rules, we considered 3 independent trials and averaged the results (see details in Appendix A.6.3).
@@ 5.2 General Continued Pre-training
We selected 20K samples from our data source using the methods described above for continued pre-training of Pythia-1B, then benchmarked the model’s performance. The choice of 20K samples was constrained by our GPU resources. Despite 20K being significantly smaller than the pre-training corpus size, we still observed improvements in benchmark results shown in Table 1, with DPP leading in most metrics. We anticipate these differences will be more obvious in domain-specific fine-tuning settings, demonstrated in 5.3 below.
@@
GPT-Uncorrelated 59.6 24.9 53.2 25.8 49.4 42.6
All 50 Rules 60 25.3 53 26.1 49.3 42.7
Random 10 Rules 60 25.1 53.7 25.8 49.5 42.8
DPP 10 Rules 60 25.7 54 26.2 50.1 43.2
Table 1: General continued pre-training of Pythia-1B using 20K selected data samples from SlimPajama (bold text denotes the first place and underlined text denotes the second place). The first row shows the original model’s performance without further training.
@@ 5.3 Domain-specifc Fine-tuning
We now focus on domain-specific fine-tuning across four domains: IMDB, Medical, Math, and Code. By selecting 20K domain-related data samples from our source for model training, we aim to enhance domain-specific task performance. As demonstrated in Tables 2 and 3, domain-specific fine-tuning yields more significant improvements than general continued pre-training, where the latter often needs larger datasets to enhance performance due to the broader nature of the training data. Notably, rule-based methods consistently outperform rule-free approaches in general, especially when comparing against Random Select and No Rule. Among all rule-based methods, QuRating underperforms. As previously noted, such human-designed rules are inherently limited due to varying preferences of designers, introducing bias in rules. Additionally, human designers may not capture the data-dependent correlation between rules effectively. The GPT-Uncorrelated rules face a similar issue where the rule selection process is entirely independent of the data. In contrast, our framework begins by automatically generating a diverse set of rules and then selecting an orthogonal subset. Furthermore, our method employs the data-dependent score vector to represent rules and utilizes a quantitative measure to accurately assess their correlation.
Within our framework, DPP demonstrated superior performance compared to using all 50 rules or selecting 10 rules randomly. This aligns with our argument of the importance of rule orthogonality, as well as the intuition that the optimal $r$ is not near the boundaries (both validated by the previous experiments on Section 4). This underscores the effectiveness of a rule-based strategy, which introduces more balanced diversity in the data rating aspects and selects better training data. Furthermore, it also demonstrates that our application of DPP in rule selection effectively identifies a core set of high-quality rules, thereby enhancing data quality and ultimately improving model performance.
@@
QuRating 47.7 21.3 42.2 22 28.5
GPT-Uncorrelated 50.9 23.7 42 22.7 29.4
All 50 Rules 51 23.1 42.6 23 29.6
Random 10 Rules 51.7 24 41.2 23.5 29.6
DPP 10 Rules 53.5 24.6 43.3 26.8 31.6
@@
Llama3-8B 41 39.6 34 38.2 46.3 42.9 44 48.4 45.4
Uniform Sampling 40.5 39.2 35 38.2 38.7 38.2 38.2 39.7 38.7
No Rule 42.3 37.4 37 38.9 45.1 43.9 42.8 52.1 45.9
DSIR 41.5 41.1 34 38.9 45.1 43.6 49.1 52.2 47.5
QuRating 41.5 38.1 35 38.2 43.2 43.4 40.5 45.6 43.1
GPT-Uncorrelated 41.4 39 37.3 39.2 41.2 43.5 39.6 48.6 43.2
All 50 Rules 41.8 40.7 33 38.5 43.9 43.4 46.6 49.1 45.8
Random 10 Rules 42.9 39.8 35.2 39.3 48.5 41 46.6 48.1 46
DPP 10 Rules 43.7 40.6 38 40.8 50.5 44.2 46.9 52.7 48.6
@@ A.2 DPP Sampling
Time Complexity: Finding the submatrix (subset of columns) of a matrix to maximize the orthogonality is NP-hard (Civril & Magdon-Ismail, 2007; Kulesza et al., 2012). DPP provides us a relatively good solution. In practice, the computational complexity of sampling from a DPP depends primarily on the eigendecomposition of the kernel matrix ${\bm{K}}$. In our case, ${\bm{K}}\in\mathbb{R}^{R\times R}$ and therefore it requires $O(R^{3})$ time, where $R$ is the number of rules. In the DPPy package (Gautier et al., 2019) it uses the spectral sampler by default, so the actual run-time of our DPP implementation is $O(R^{3})$.
DPP Sampling for Data Selection: We noticed that in a concurrent work Yang et al. (2024), the authors also use DPP to perform data selection, but directly applied to the data itself. However, the approach to directly perform data selection using DPP requires the computation based on the kernel matrix with dimension $N$ (number of samples), which is usually huge in the context of LLM data. Moreover, while DPP inherently prioritizes diversity in data selection, it does not address other quality dimensions. In contrast, our rule-based approach assesses multiple aspects of data quality, ensuring a more comprehensive and robust selection process.
@@ A.5.4 Rater Model Size Variation: Llama3-70B-Instruct
A.5.5 Rule Generator Variation: Claude-3.5-Sonnet
To verify that GPT-4 is a reliable rule generator, we compare it with Claude-3.5-Sonnet. For each of the five tasks (General, IMDB, Medical, Math, Code), we prompt GPT and Claude to generate 100 rules for each, and then study the distribution of the rules. Specifically, we use Sentence-Transformer (Reimers, 2019) to generate the embedding vectors and then use PCA to project them onto the top two principal components for 2-dimensional visualization. From Figure 8 below, we observe that the two groups of rules generated by the two models completely overlap, demonstrating no distinct separation. This suggests that GPT-4 functions effectively as an unbiased rule generator.
(a) General Rules
(b) IMDB Rules
(c) Medical Rules
(d) Math Rules
(e) Code Rules
Figure 8: Embeddings of rules generated by GPT and Claude across different domains.
To further quantify the distribution differences, we studied the Wasserstein distance within and between the two rule sets. Specifically, we compute the distance between the GPT rules and Claude rules. Then we randomly split GPT rules into two parts and computed their Wasserstein distance and similarly for the Claude rules (averaged over 10 trials). By comparing these values (see Table 4), we found no clear distribution bias when switching from one rule generator to the other.
Table 4: Comparison of Intra-Model and Inter-Model Metrics across Domains
General IMDB Medical Math Code
Intra-GPT 0.432 0.438 0.502 0.584 0.673
Intra-Claude 0.455 0.445 0.515 0.592 0.668
Inter-Model 0.468 0.548 0.522 0.612 0.684
Now we have compared the rule generators using GPT and Claude above. In order to address potential biases in rule generation by these large language models (LLMs) compared to human-generated rules, we prompt GPT-4 to generate 133 ethical and safety rules and compare the GPT-generated rules with the public and standard constitutions in Huang et al. (2024) (we make the rules all start from “Choose the response that” for a fair comparison). We asked 3 authors who have not seen the constitutions in Huang et al. (2024) to distinguish the rules blindly. We get an average accuracy of $20.7\%$, suggesting it is indeed hard to distinguish between rules generated by GPT and those designed by humans. All these discussions underscores the potential of GPT as a reliable rule generator that is capable of producing rules that are comparable to those crafted by human experts.
A.5.6 Prompts and Generated Rules